A 2 kg ball moving at 2 m/s collides with a stationary 1 kg ball. What is the speed of the 1 kg ball after the collision if the collision is:

(a) elastic, and the speed of the 2 kg ball drops to 1 m/s?

(b) inelastic?
Answer #1

a)

if the collision is elastic

v1 = 2 m1 u1/ ( m1 + m2)

v1 = 2* 2* 2 / ( 2 + 1) = 2.667 m/s

======

b)

if collision in inelastic

m1 u1 + m2 u2 = m1 v1 + m2 v2

2 * 2 + 0 = 2 * 1 + 1* v

v = 2 m/s

======

Comment before rate in case any doubt, will reply for sure.. goodluck